# Premerger phenomena in neutron-star binary coalescences

### v3

@@ VII.1 A brief look at future possibilities
One of the key ways in which multimessenger astrophysics will be propelled forward in the future comes from next-generation GW interferometers. Although there are a number of proposed technologies, some of the most notable are (i) the Einstein Telescope (ET; [263, 335, 190, 67]). The sensitivity band of ET is similar to that of the LVK network, though notably deeply with considerable improvement in the $\lesssim 10$ Hz range. The reason is that the gravitational gradient (aka “Newtonian”), seismic, and thermal noises are less of a problem for ET, as it will be built underground and include cryogenic technologies to reduce thermal vibrations [335]. This latter band is relevant to capture the onset of low-frequency oscillations and track the orbit for longer periods of time. (ii) Cosmic Explorer (CE; [6, 348, 135]). This detector is similar to ET in terms of sensitivity band, though is planned to go even deeper. CE may detect thousands of neutron stars in merger [136], and thus lead to a true era where statistical methods can nail down the neutron-star EOS. (iii) The Laser Interferometer Space Antenna (LISA). LISA is a planned space-based detector that could capture the very early aspects of inspiral while the orbit is at $\gtrsim$ mHz frequencies [21]. Wide binary elements could be useful for ocean resonances [387, 386], low frequency $g$-modes (assuming these are not washed out by reaction rates; Sec. III.3) and the testing of spacetime structure generally [see, e.g. 115, 116].
@@
Innovations on the electromagnetic side are no less promising. For instance, with the development of the SKA and next-generation X-ray observatories the types of premerger precursors could increase significantly. Although we have limited our use of the word precursor to specifically mean the gamma-ray flashes occurring before the main event (see Tab. 3), it is hoped that such terminology will prove ambiguous in the future. Tables 1 and 2 of [177] provide a summary of next-generation telescopes that could coincidentally detect signals associated with neutrino bursts originating from a binary merger. One intriguing possibility comes from the planned POLAR-2 detector, which should able to measure polarization and similar effects related to the propagation of gamma-rays [a survey of other next-generation gamma-ray facilities are described in Section 2 of 202]. This could be used to probe the local magnetic field strength in the region where the rays were produced. This would allow one to test, for instance, the extent to which non-thermality is tied to magnetars in mergers (see also Sec. IV.4).

### v4

@@ VII.1 A brief look at future possibilities
One of the key ways in which multimessenger astrophysics will be propelled forward in the future comes from next-generation GW interferometers. Although there are a number of proposed technologies, some of the most notable are (i) the Einstein Telescope (ET; [263, 335, 190, 67]). The sensitivity band of ET is similar to that of the LVK network, though notably deeply with considerable improvement in the $\lesssim 10$ Hz range. The reason is that the gravitational gradient (aka “Newtonian”), seismic, and thermal noises are less of a problem for ET, as it will be built underground and include cryogenic technologies to reduce thermal vibrations [335]. This latter band is relevant to capture the onset of low-frequency oscillations and track the orbit for longer periods of time. (ii) Cosmic Explorer [CE; 6, 348, 135]. This detector is similar to ET in terms of sensitivity band, though is planned to go even deeper. CE may detect thousands of neutron stars in merger [136], and thus lead to a true era where statistical methods can nail down the neutron-star EOS. (iii) The Laser Interferometer Space Antenna (LISA). LISA is a planned space-based detector that could capture the very early aspects of inspiral while the orbit is at $\gtrsim$ mHz frequencies [21]. Wide binary elements could be useful for ocean resonances [387, 386], low frequency $g$-modes (assuming these are not washed out by reaction rates; Sec. III.3) and the testing of spacetime structure generally [see, e.g. 115, 116].
@@
Innovations on the electromagnetic side are no less promising. For instance, with the development of the SKA and next-generation X-ray observatories the types of premerger precursors could increase significantly. Although we have limited our use of the word precursor to specifically mean the gamma-ray flashes occurring before the main event (see Tab. 3), it is hoped that such terminology will prove ambiguous in the future. Tables 1 and 2 of Guépin et al. 177 provide a summary of next-generation telescopes that could coincidentally detect signals associated with neutrino bursts originating from a binary merger. One intriguing possibility comes from the planned POLAR-2 detector, which should able to measure polarization and similar effects related to the propagation of gamma-rays [a survey of other next-generation gamma-ray facilities are described in Section 2 of 202]. This could be used to probe the local magnetic field strength in the region where the rays were produced. This would allow one to test, for instance, the extent to which non-thermality is tied to magnetars in mergers (see also Sec. IV.4).